OpenStudy (maheshmeghwal9):

Two conducting spheres P & Q have radii 'R' & 'r' respectively. They are connected by a thin wire, charged and then separated to a large distance. Now at the surface of P & Q, the ratio of electric fields is: - A.) \(\Large{r/R}\) B.) \(\Large{R^2/r^2}\) C.) \(\Large{r^2/R^2}\) D.) None of these.

OpenStudy (maheshmeghwal9):

@experimentX plz help:)

OpenStudy (maheshmeghwal9):

@Rohangrr @ajprincess @UnkleRhaukus Plz help :)

OpenStudy (anonymous):

none of these

OpenStudy (maheshmeghwal9):

but in my book answer is A.) r/R

OpenStudy (anonymous):

because it says " P & Q have radii 'R' & 'r' respectively. They are connected by a thin wire, charged and then separated to a large distance. Now at the surface of P & Q, the ratio of electric fields is" it should be

OpenStudy (maheshmeghwal9):

but my book's answer is different:\

OpenStudy (maheshmeghwal9):

it is option A.

OpenStudy (anonymous):

correct I didnt look the last bit of it

OpenStudy (anonymous):



OpenStudy (experimentx):

when two charges spheres are connected by a thin wire ... the whole potential of both spheres must be equal.

OpenStudy (maheshmeghwal9):

i haven't studied potential yet:(

OpenStudy (maheshmeghwal9):

is there any solution to this question based on gauss's law?

OpenStudy (maheshmeghwal9):

becoz i have studied \[E*A=\frac{q}{\epsilon_0}.\]

OpenStudy (maheshmeghwal9):

or any other solution without the concept of electric potential?

OpenStudy (experimentx):

thinking

OpenStudy (maheshmeghwal9):

ok:)

OpenStudy (maheshmeghwal9):

I have found this much \[E_1=\frac{q}{4 \pi R^2 \epsilon_0}\] \[E_2=\frac{q}{4 \pi r^2 \epsilon_0}\] \[\frac{E_1}{E_2}=\frac{r^2}{R^2}.\]q = charge gained by P = charge lost by Q but this is wrong; i know:(

OpenStudy (experimentx):

probably they have charge distribution according to their sizes

OpenStudy (maheshmeghwal9):

may be.

OpenStudy (experimentx):

since it's sphere ... it must be symmetric. \[ E_2=\frac{(r/R)q}{4 \pi r^2 \epsilon_0}\] don't put those 4 pi r^2 unless it's necessary

OpenStudy (maheshmeghwal9):

sorry i don't get u how u gt E_2 as above?

OpenStudy (maheshmeghwal9):

i mean from where did this r/R come?

OpenStudy (experimentx):

i told you .. both don't have same charges.

OpenStudy (maheshmeghwal9):

but if they are connected through a wire they must gain or lose equal amount of charges i think

OpenStudy (experimentx):

R has Rq charge and r as rq charge. depending on radii

OpenStudy (experimentx):

if R as q charge then r as r/R q charge.

OpenStudy (experimentx):

going offline ... i'll won't be back in less than 2 hrs

OpenStudy (maheshmeghwal9):

np:)

OpenStudy (shane_b):

r/R makes perfect sense because the surface area of a sphere is:\[SA_{sphere}=4\pi r^2\]Conducting spheres will hold charges proportional to their surface area. You don't need Gauss' Law for this.

OpenStudy (maheshmeghwal9):

but how to get r/R mathematically?

OpenStudy (shane_b):

\[\frac{4pir^2}{4piR^2} \propto \frac{r}{R}\]

OpenStudy (shane_b):

Actually, I guess you'd really get:\[\frac{4pir^2}{4piR^2} = \frac{r^2}{R^2}\]

OpenStudy (maheshmeghwal9):

really confusing:\

OpenStudy (shane_b):

The more I think about it the more I convince myself that regardless of what the book says, the answer should be r^2/R^2 :/

OpenStudy (maheshmeghwal9):

i also thought this first time as in my solution:) i think we should wait for others also to confirm.

OpenStudy (shane_b):

Yes :)

OpenStudy (maheshmeghwal9):

:)

OpenStudy (shane_b):

I found on an old exam online which almost makes sense: Since the spheres are connected by a conducting wire, they are initially at the same potential:\[k\frac{q_1}{r_1} = k\frac{q_2}{r_2} => \cancel{k}\frac{q_1}{q_2} = \cancel{k}\frac{r_1}{r_2}\]The electric field at the surface of each shell is: \[k\frac{q}{r^2}\]Thus the ratio of electric fields is:\[\frac{E_1}{E_2} = (\frac{q_1}{q_2})(\frac{r_2^2}{r_1^2})\]From the previous equality:\[\frac{E_1}{E_2} = \frac{r_2}{r_1}\]

OpenStudy (maheshmeghwal9):

thank u very much:D

OpenStudy (shane_b):

Maybe I'm just not awake yet but I'm still not seeing how we go from r^2/R^2 to r/R....unless the spheres are assumed to have the same radius in which case the question would not have listed r and R.